How many moles of air are contained In the lungs of an average healthy adult with a total lung capacity of 3.8L?
Answer #1

Usin gthe ideal gas equation: n = PV / RT, P is pressure, V is volume, R is gas constant and T is temperature in K. The temperature is assumed to be room temperature and pressue

n = 1 * 3.8 / (0.0821 * 298) = 0.155 moles